Write a short paragraph of about ten lines about your experience of learning how to work with fractions
1
Expert's answer
2020-09-22T13:45:17-0400

Mostly fractions are always expressed in their simplest forms, for example "\frac{4}{16}" has both the numerator and denominator divisible by 4 thus it can be expressed in it's simplest form as "\frac{1} {4}". The simplest form of a fraction means that both the numerator and denominator have no common divisor or factors.Fractions can also represent quantities greater than one for example"\frac{7}{3 },\frac{9}{4}" . Fractions represents a number but also a ratio. Fractions need to belong to the same ‘family’ in order to add or subtract them meaning the denominator for the fractions to be added or subtracted must be the same.In other cases, this requires adjusting the fractions so that they

have a common denominator. It is important that this adjustment preserves the ratio between the numerator and the denominator. This adjustment to the same ‘family’ is not necessary when multiplying or dividing fractions.